please help me solve by elimination
5x+6y=3
10x+12y=6
answer is ordered pair integers or fractions

Found 2 solutions by Alan3354, drj:
Answer by Alan3354(21608)

Eqn 2 is 2x Eqn 1 --> dependent equations.
No unique solution is possible.

Answer by drj(1380)

Phelp me solve by elimination

5x+6y=3 (Equation 1) or multiplying by 2 to both sides of the equation yields 10x+12y=6
10x+12y=6 (Equation 2)

Answer is ordered pair integers or fractions

Equation 1 is identical to Equation 2 if you multiply Equation 1 by 2. Therefore you have an infinite set of solutions. When I do multiply Equation 1 by 2 and subtract Equation 2 from multiplied version of Equation 1, I get 0=0 on both sides of the equation while eliminating both the x and y terms.

ANSWER: There are an infinite set of solutions satisfying the system of equations.

I hope the above steps were helpful.

For free Step-By-Step Videos on Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra or for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J

http://www.FreedomUniversity.TV
http://www.Facebook.com/FreedomUniversityTV
http://www.Twitter.com/FreedomUTV